# Quasiperiodicity protects quantized transport in disordered systems without gaps

### v1

@@ III Robustness to spatial disorder
We numerically compute the pumped charge as the average number of sites traveled per pump cycle from a linear fit of the center of mass motion for a single particle initially localized in a band with Chern number $C=+1$ (Fig. 3 a) and $+4$ (Fig. 3 f). For small disorders (blue shaded area), the extracted values agree as expected with the above quantized predictions in the presence of global gaps, see Fig. 3 e, j. However, the quantized currents survive far beyond the closing of the clean energy gaps (dashed blue lines), and break down only at larger critical disorder strengths (dashed black vertical lines). In the $C=+4$ band shown in Fig. 3 j, we find that the critical disorder strength for this particular band exceeds the size of the clean $\Delta E^{(\pm 2)}$ gaps by more than an order of magnitude. These critical disorder strengths are band-dependent even if the bands share a gap, see Appendix B.
Once more, configuration space provides a simple local picture for the stability of the quantized current as well as the band-dependent critical disorder strength. In the clean case, the phase winding $\Delta\varphi$ between two consecutive Landau-Zener processes for a given localized state is constant (Fig. 4 a,b) and corresponds to the angle spanned by the corresponding resonance on the unit circle, see Fig. 2d. In the presence of disorder, this phase fluctuates randomly, giving rise to a distribution of spatially-dependent and randomly distributed $\Delta\varphi_{i}=\Delta\varphi+\delta\varphi_{i}$. This is reflected in the fluctuating size of the plateaus in Fig. 3 c. As long as the disorder is below a critical value (which we calculate in Appendix B), these fluctuations $\delta\varphi_{i}$ average to zero and leave the quantized current unaffected (Fig. 4 c,d). Once the disorder becomes strong enough to shrink one ore more of the $\Delta\varphi_{i}$ down to 00 (Fig. 4 e,f), the order of resonances may change, which results in the breakdown of the quantized current as the particle starts being back-reflected.
Thus, the robustness of the quantized current is not limited by the size of the clean energy gaps, but instead depends on the ordering of consecutive local resonances. In the clean case, these resonances are all aligned in energy, giving rise to a global gaps. Increasing onsite disorder shifts each resonance independently and thereby closes the global energy gaps, but does not necessarily modify their ordering. As long as the order between consecutive local resonances remains unaffected, quantized currents can subsist, even in the complete absence of global energy gaps in the system. Notably, bands for which $\Delta\varphi$ is larger, i.e., those which have a larger IDoS, are more robust to fluctuations $\delta\varphi_{i}$. This constitutes a novel form of topological protection.
@@
Figure 4: Local argument for the stability and breakdown of quantized currents: Coloured lines in the left column panels are on-site energies of consecutive neighbors. Thick coloured line illustrates the adiabatic trajectory of the state under pumping for $V\gg J$, whose real-space density is shown in the right column. (a,b) In the clean case, the phase $\Delta\varphi$ separating two consecutive Landau-Zener transitions is constant, and the particle undergoes successive adiabatic transitions to neighboring sites at a constant rate (blue curve). (c,d) The addition of not-too-large onsite disorder induces fluctuations in the phase $\Delta\varphi$ between transitions, but leaves the pumped current quantized. (e,f) When the disorder strength exceeds a critical value, the order of successive Landau-Zener processes can be changed, causing the particle to be backreflected. The argument is shown here for the top $C=+1$ first-order band, but can be easily applied to any other band. Parameters: $\beta=\sqrt{2}/2$.
@@ IV Experimental Signatures and Quantum State Preparation
The above procedure generates a 1D many-body state with an associated virtual Chern number. To generate a state with a well-defined “real” 2D Chern number, the protocol can in principle be extended to a 2D lattice consisting of an array of $n_{y}$ AAH chains with increasing phase $\varphi(y)=2\pi y/n_{y}$ along a second spatial dimension $y$. This results in a 2D Hamiltonian given by
$\hat{H}_{2D}=\sum_{(x,y)}$ $2V\cos\left(2\pi\beta x+\frac{2\pi}{n_{y}}y\right)\hat{a}_{x,y}^{\dagger}\hat{a}_{x,y}$
$-J\left(\hat{a}_{x+1,y}^{\dagger}\hat{a}_{x,y}+\hat{a}_{x,y}^{\dagger}\hat{a}_{x+1,y}\right)\,,$ (18)
whose spectrum is the same as that of the 1D AAH Hamiltonian [46, 34]. Under Fourier transforming $y\rightarrow\tilde{y}$, Section IV is equivalent to the 2D Harper Hamiltonian
$\hat{H}_{QHE}=\sum_{(x,\tilde{y})}$ $V\left(e^{i2\pi\beta x}\hat{a}_{x,\tilde{y}+1}^{\dagger}\hat{a}_{x,\tilde{y}}+h.c.\right)$
$-J\left(\hat{a}_{x+1,\tilde{y}}^{\dagger}\hat{a}_{x,\tilde{y}}+h.c.\right),$ (19)
which describes a square lattice with magnetic flux $2\pi\beta$ per plaquette. Therefore, by stacking multiple copies of the one-dimensional state preparation protocol, we prepare a state with a desired Chern number of the 2D Harper Hamiltonian, albeit in a mixed real-momentum space configuration. The chains can be coupled by introducing small tunneling terms between neighboring chains without altering the Chern number of the state, provided that the spectral gaps remain open [47].
Interestingly, on-site interactions effectively appear as local momentum-space interactions in the $y$ coordinate [48]. This mixed real-momentum space system presents an opportunity to probe new quantum Hall physics inaccessible to traditional condensed-matter systems, e.g. one could engineer local scalar potentials in momentum space to reproduce dispersion relations corresponding to long-range real-space hopping models which have been used to generate exact fractional quantum hall (FQH) phases on a lattice via topological flat bands [49], and non-local real-space interactions have been used to stabilize non-abelian fractional Chern insulator (FCI) phases in numerics [50]. Moreover, numerical studies of fractional Chern insulators have long noticed enhanced stability of the FCI phase in higher Chern number bands [51] along with new phases of matter with no FQH analogs [52]. As such, the above state preparation protocol could become a new experimental platform for exploring numerous theoretical proposals involving large Chern numbers, flat bands, and non-local interactions.
@@ V Discussion and Conclusion
We present a novel example of topological protection in quasiperiodic systems depending on the number of states and the position of the band in configuration space rather than a spectral gap condition, challenging the traditional understanding that the robustness of topological phenomena depends on the presence of spectral gaps [37], mobility gaps [17], or symmetry protections [19]. By tracking the fate of quantized currents in the disordered Aubry-André chain (1D almost Mathieu operator), we find that the combination of quasiperiodic Anderson localization and the gap labelling theorem results in a robustness to bounded local disorder far beyond the closing of the relevant spectral gaps. We furthermore develop a local picture to explain quasiperiodic Thouless pumping both in the clean and disordered case that provides excellent estimates for the numerically observed critical disorder strengths.
@@
The outlined mathematical relations to ergodicity and the Dry Ten Martini problem [26, 29, 28] ensure pathways to generalizations. In one-dimension, it would be worth exploring the case of several incommensurate frequencies, which are likely to result in enhanced stabilization mechanisms that in turn should lead to novel measurable signatures. These pursuits can also be followed in the context of the Dry Ten Martini problem without dualities or with hidden dualities [29, 28, 53], where localization phase transitions may be richer.
@@
In two-dimensional systems, the presence of second Chern numbers and a richer set of topological invariants could lead to more examples of gap-independent topological protection. For example, in ultracold atomic systems, the scheme could be extended to the two-dimensional Aubry-André model or optical quasicrystals [41, 42, 43], where pumping could be realised along multiple dimensions. The recent advent of moiré van der Waals structures, in which pumping can be achieved upon applying periodic external gauge fields, provides for an exciting solid-state platform in which these new mechanisms might both occur and help explain experimental observations.
@@
In summary, our results uncover a novel protection mechanism for Thouless pumping, highlighting an unexplored richness of topology in quasiperiodic systems and motivating a re-examination of topological protection in the absence of translation invariance. They furthermore enable novel experimental applications, notably in the preparation of quantum states with high Chern numbers.
@@ VI Acknowledgements
We cordially thank Joel E. Moore for insightful feedback and advice regarding the draft. We also thank Arjun Ashoka, Alexander Avdoshkin, Vir B. Bulchandani, Fabian Heidrich-Meisner, Matthew Henricks, Johannes Mitscherling, Suman Mondal, Lee Reeve, Ruben Verresen, Qijun Wu, and Maciej Zworski for many helpful discussions. R.-J.S. acknowledges funding from a New Investigator Award, EPSRC grant EP/W00187X/1, a EPSRC ERC underwrite grant EP/X025829/1, and a Royal Society exchange grant IES/R1/221060 as well as Trinity College, Cambridge. Work by D.S.B. was performed with support through a MURI project supported by the Air Force Office of Scientific Research (AFSOR) under grant number FA9550-22-1-027. U.S. and E.G. acknowledges support by the European Commission ERC Starting Grant QUASICRYSTAL, the EPSRC Grant (No. EP/R044627/1), and EPSRC Programme Grant DesOEQ (No. EP/P009565/1). E.G. acknowledges support from the Cambridge Trust. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ VII Authors contributions
Appendix A Gap labels from perturbation theory
@@
We showed in Section I how resonant pairs of lattice sites coupled by the effective $n$-th order tunneling element $J_{eff}^{(n)}$ gave rise to pairs of energy gaps in the spectrum of the AAH chain at energies ±$\pm 2V\cos{(n\pi\beta)}$ ). According to the gap labelling theorem, each of these gaps is characterised by a Chern number, whose sign dictates the direction of the current arising under adiabatic pumping when all states below the gap are occupied.
@@
In the case $[n\beta]$ even, let us for instance fill up all states up to the energy $-2V\cos{(n\pi\beta)}$ ), see Fig. 7. We note that for the case discussed in Fig. 2 ($n=1$, $\beta>0.5$), this energy is positive. For states below this gap, the pumping induces Landau-Zener hoppings in the negative direction, where each Landau-Zener transition transfers the states from the right-hand resonant neighbor to the left-hand one (we can ignore the Landau-Zener processes associated with the gaps below $-2V\cos{(n\pi\beta)}$ ), as they do not influence the overall current). This results in a quantized current in the negative direction and this gap is therefore associated with a negative Chern number. The same reasoning can be applied to the state filled up to $2V\cos{(n\pi\beta)}$ ), leading to a positive gap label.
@@
In the case $[n\beta]$ odd, the resonant neighbors are flipped, and the adiabatic Landau-Zener transition transfers the states from the left- to the right-hand neighbor. In this case, the quantized current flows in the positive direction and is therefore associated to a positive Chern number.
@@
Therefore, to assign the correct sign to the labels of the energy gaps, one has to treat both the cases of even and odd parity of $n\beta$, as done in Eq. 9.
@@
Figure 7: Dependence of the signs of the gap labels on the parity of $[n\beta]$: (a) When $[n\beta]$ is even, the associated Landau-Zener transition involves transport from the right-hand resonant neighbor to the left-hand one. The resulting quantized current flows in the negative direction is therefore associated with a negative Chern number. (b) When $[n\beta]$ is odd, the situation is reversed, and the quantized current, which flows in the positive direction, must be associated with a positive Chern number. Note that in contrast to Fig. 2, we are here considering states below the gap.
@@
Appendix B Critical disorder strengths for first-order bands

### v3

@@ III Robustness to spatial disorder
We numerically compute the pumped charge as the average displacement, that is the average number of sites traveled, per pump cycle from a linear fit to the center of mass motion for a single particle initially localized at a configuration-space position which corresponds to a band with Chern number $C=+1$ (Fig. 3 a) and $+4$ (Fig. 3 f) in the clean case $W=0$. For small disorders below the closing of the relevant gaps (blue shaded area), the extracted values agree as expected with the above quantized predictions in the presence of global gaps, see Fig. 3 e, j. However, the quantized currents survive far beyond the closing of the clean energy gaps (dashed blue lines), and break down only at larger critical disorder strengths. In the $C=+4$ band shown in Fig. 3 j, we find that the critical disorder strength for this particular band exceeds the size of the clean $\Delta E^{(\pm 2)}$ gaps by more than an order of magnitude.
@@
Figure 4: Local picture for the stability and breakdown of quantized currents: Coloured lines in the left column panels denote onsite energies of three consecutive sites for $V\gg J$. Thick coloured line illustrates the adiabatic trajectory of the state under pumping. The right column shows the corresponding real-space density. (a,b) In the clean case, the phase $\Delta\varphi$ separating two consecutive Landau-Zener transitions is the same for each transition and the particle undergoes successive adiabatic transitions to neighboring sites at a constant rate (blue curve). (c,d) The addition of not-too-large onsite disorder induces fluctuations in the phase $\Delta\varphi$ between transitions but preserves the order of resonances and leaves the pumped current quantized. (e,f) When the disorder strength exceeds a critical value, the order of successive Landau-Zener processes can change, causing the particle to be back-reflected. The argument is shown here for the top $C=+1$ first-order band. Parameters: $\beta=\sqrt{2}/2$.
Once more, configuration space provides a simple picture for the stability of the quantized current. In the clean case, the phase winding $\Delta\varphi$ between two consecutive Landau-Zener processes for a given localized state is constant (Fig. 4 a,b) and corresponds to the angle spanned by the corresponding band on the unit circle, see Fig. 2d. In the presence of disorder, this phase fluctuates randomly, giving rise to a distribution of spatially-dependent and randomly distributed $\Delta\varphi_{i}=\Delta\varphi+\delta\varphi_{i}$. This is reflected in the fluctuating size of the plateaus in Fig. 3 c. As long as the disorder is below a critical value, these fluctuations $\delta\varphi_{i}$ average to zero and leave the quantized current unaffected (Fig. 4 c,d). Once the disorder becomes strong enough to shrink one ore more of the $\Delta\varphi_{i}$ down to 00 (Fig. 4 e,f), the order of resonances may change, which results in the breakdown of the quantized current as the particle starts being back-reflected.
The above analysis rests on the assumption that the pumping rate is high enough to prevent the occurrence of Landau-Zener tunneling over longer ranges than desired. To illustrate this necessary condition, we investigate the impact of the pumping rate on the critical disorder strength for destroying the quantized currents in the $C=+1$ bands. Fig. 5 shows the relative error in the quantization of the displacement per cycle of the right-moving part of an initial state consisting of a product of $20$ local Fock states on neighboring sites, see Fig. 6 for reference. The red dotted line shows the configuration-space estimate
$W_{crit}^{C=+1}=2V\sin^{2}{\left(\pi\beta\right)}.$ (18)
of the disorder strength necessary for the condition $\Delta\varphi_{i}=0$, see details in Appendix B.
The quantized current emerges for pumping rates low enough to be adiabatic with respect to the largest spectral gaps of order $J$, i.e., for $\dot{\varphi}\approx\mathcal{O}(2J^{2}/V)$. The quantization breaks down again when the pumping becomes adiabatic with respect to the next spectral gap of order $J_{\textit{eff}}^{(2)}$, i.e., $\dot{\varphi}\approx\mathcal{O}(2J^{4}/V^{3})$, and hence enables undesired longer-range tunneling. Consistent with [15], the protection of the quantization to disorder closing spectral gaps is not expected to hold in the full adiabatic limit, where Landau-Zener transitions can take place over arbitrary distances. Interestingly, we observe that for slow pumping the addition of weak disorder induces a stabilization of the quantization, owing to the suppression of pumping via longer-range resonances. The critical disorder strength observed in Fig. 5, where the initial state is a product state of many particles, is in good agreement with the single-particle results shown in Fig. 3 e.
Figure 5: Effect of pumping rate on the quantization of the current in the $C=+1$ band: Color denotes the average relative error in the quantization of the displacement per cycle (extracted from the slope of the right-moving part of the density $|\Psi(x>0)|^{2}$ of an initial state consisting of a product of $20$ local Fock states on neighboring sites, cf. Fig. 6). Black contour lines indicate the regions with less than $5\%$ error. Blue dotted line indicates the closing of the first-order energy gaps due to disorder. Red dotted line indicates the configuration-space estimate for the critical disorder strength (Eq. 18). For very low pumping rates, current stops being quantized once unwanted longer-range tunneling processes become relevant. These additionally increase the sensitivity of the pump to the disorder. The increasing separation of timescales between the first- and second-order gaps for larger $V/J$ allows for a better protection of the quantization for large $V/J$. Parameters: (a) $V/J=7.5$. (b) $V/J=15.0$. (c) $V/J=30.0$. Pumped angle is $15\pi$. Each point is obtained after averaging over $20$ disorder realizations.
Thus, provided that the pumping rate is set correctly to suppress undesired longer-range tunneling processes, the robustness of the quantized current is not limited by the size of the clean energy gaps, but instead depends on the ordering of consecutive local resonances. In the clean case, these resonances are all aligned in energy, giving rise to global gaps. Increasing onsite disorder shifts each resonance independently and thereby closes the global energy gaps, but does not necessarily modify their ordering. As long as the order between consecutive local resonances remains unaffected, quantized currents can subsist, even in the complete absence of global energy gaps in the system.
The above robustness is distinct from the situation in conventional disordered topological systems studied in the context of the integer and fractional quantum hall effects [18], where the robustness beyond the closing of the spectral gap stems from the persistence of a mobility gap between the (gapped) topological extended and (gapless) localized states of the system [43] and is often described using real-space topological invariants called Chern markers [44]. However, this picture does not directly apply to the 1D AAH Hamiltonian as it only hosts insulating states for $V>J$. At first glance, the notion of extended states may be restored by treating $\varphi$ as a virtual dimension in $\hat{H}^{J,V,W}(\varphi)$ and applying the famous mapping of the 1D AAH Hamiltonian to the 2D Hofstadter Hamiltonian [35], see Section IV. However, the onsite disorder $\hat{W}$ considered here only generates a striped disorder pattern in 2D that cannot localize states along the virtual dimension [45]. Thus, all states are extended and the traditional mobility gap closes.
Interestingly, the Landau-Zener formalism above implicitly depends on the finite pumping period $T$ to limit long-range resonances in the presence of strong disorder [15], see Fig. 5. This period also appears in the 1D to 2D mapping when considering the full Schrödinger operator, $\hat{K}\equiv i\partial_{t}-\hat{H}^{J,V,W}(\varphi(t))$. Since $\hat{H}^{J,V,W}(\varphi(t))$ is periodic in time with period $T$, only discrete frequencies $\omega=\dot{\varphi}\tilde{j}$ with $\tilde{j}\in\mathbb{Z}$ appear when Fourier transforming $\hat{K}$ along the time coordinate, forming a discrete virtual dimension [7, 46]. The Fourier transform furthermore exchanges the time-derivative for a linear potential proportional to the pumping rate $\dot{\varphi}$, $\tilde{K}\equiv\sum_{\tilde{j}\in\mathbb{Z}}\dot{\varphi}\tilde{j}-\hat{H}_{QHE}-\hat{W}\otimes\mathbb{I}_{\tilde{j}}$ [47], see Section IV for definition of $\hat{H}_{QHE}$. The eigenfunctions of the Schrödinger operator, i.e. those states satisfying $\tilde{K}\psi_{\lambda}=\lambda\psi_{\lambda}$, form a diagonal basis for the unitary $\hat{U}(T)$ describing the stroboscopic time evolution of states in the 1D Thouless pump. Thus, a possible reconciliation is that the finite pumping rate introduces a non-vanishing electric field (the linear potential) along the Fourier-transformed virtual $\tilde{j}$ direction and induces Wannier-Stark localization along this direction [48]. The mix of Wannier-Stark and Anderson localization localizes topologically trivial states and generates a hybrid mobility gap. This suggests the existence of finer definitions for mobility gaps in the Thouless pumping of quasiperiodic systems. Determining whether or not these two pictures are indeed equivalent constitutes a promising direction for future work.
@@ IV Experimental Signatures and Quantum State Preparation
The above procedure generates a 1D many-body state with an associated virtual Chern number. To generate a state with a well-defined “real” 2D Chern number, the protocol can in principle be extended to a 2D lattice consisting of an array of $n_{y}$ AAH chains with increasing phase $\varphi(j)=2\pi j/n_{y}$ along a second spatial dimension $y$. This results in a 2D Hamiltonian given by
$\hat{H}_{2D}=\sum_{(i,j)}$ $2V\cos\left(2\pi\beta i+\frac{2\pi}{n_{y}}j\right)\hat{a}_{i,j}^{\dagger}\hat{a}_{i,j}$
$-J\left(\hat{a}_{i+1,j}^{\dagger}\hat{a}_{i,j}+\hat{a}_{i,j}^{\dagger}\hat{a}_{i+1,j}\right)\,,$ (19)
whose spectrum is the same as that of the 1D AAH Hamiltonian [54, 35]. Under Fourier transforming $j\rightarrow\tilde{j}$, Section IV is equivalent to the 2D Harper Hamiltonian
$\hat{H}_{QHE}=\sum_{(i,\tilde{j})}$ $V\left(e^{i2\pi\beta i}\hat{a}_{i,\tilde{j}+1}^{\dagger}\hat{a}_{i,\tilde{j}}+h.c.\right)$
$-J\left(\hat{a}_{i+1,\tilde{j}}^{\dagger}\hat{a}_{i,\tilde{j}}+h.c.\right),$ (20)
which describes a square lattice with magnetic flux $2\pi\beta$ per plaquette. Therefore, by stacking multiple copies of the one-dimensional state preparation protocol, we prepare a state with a desired Chern number of the 2D Harper Hamiltonian, albeit in a mixed real-momentum space configuration. The chains (momentum classes) can be coupled by introducing small tunneling elements between neighboring chains without altering the Chern number of the state, provided that the spectral gaps remain open [55].
Interestingly, onsite interactions effectively appear as local momentum-space interactions along the $\tilde{j}$ coordinate [56]. This mixed real-momentum space system presents novel opportunities to probe new quantum Hall physics inaccessible to traditional condensed-matter systems by for example engineering scalar potentials in momentum space to reproduce dispersion relations corresponding to long-range real-space hopping models that have been used to generate exact fractional quantum hall (FQH) phases on a lattice via topological flat bands [57]. Similarly, non-local real-space interactions have been used to stabilize non-abelian fractional Chern insulator (FCI) phases in numerics [58]. Moreover, numerical studies of fractional Chern insulators have long noticed enhanced stability of the FCI phase in higher Chern number bands [59] along with new phases of matter with no FQH analogs [60]. As such, the above state preparation protocol could become a new experimental platform for exploring numerous theoretical proposals involving large Chern numbers, flat bands, and non-local interactions.
@@ V Discussion and Conclusion
We present a novel example of topological protection in quasiperiodic systems that does not depend on the presence of spectral gaps [39], symmetry protections [20], or even traditional mobility gaps [18]. By tracking the fate of quantized currents in the disordered Aubry-André chain (1D almost Mathieu operator), we find that the combination of quasiperiodic Anderson localization and the gap labeling theorem results in a robustness to bounded local disorder far beyond the closing of the relevant spectral gaps. We furthermore develop a local picture to explain quasiperiodic Thouless pumping both in the clean and disordered case to explain the observed robustness.
As mentioned in the main text, determining whether or not the local picture developed here and the global breakdown of mixed linear/disorder potential mobility gaps are indeed equivalent constitutes a promising direction for future work. In particular, for weak disorder, where (large) spectral gaps of the 1D AAH Hamiltonian remain open, the infinite-range quasiperiodic resonances do not obstruct quantized pumping in low-order bands as $\dot{\varphi}\rightarrow 0$, despite the continued bifurcations into smaller bands. For strong disorder, on the other hand, finite pump frequencies play an important role in protecting the quantized currents by truncating long-range Landau-Zener transitions and possibly establishing a hybrid mobility gap, consistent with the picture in Ref. [15]. This change in the role of long-range resonances with increasing disorder suggests a potentially interesting region of parameter space for future study where the character of the 1D eigenstates is preserved, but the topological invariants of the 2D Hamiltonian are not protected by a mobility gap. Furthermore, coupling many of these 1D chains into a 2D system presents novel opportunities to probe new quantum Hall physics inaccessible to traditional condensed-matter systems.
@@
The outlined mathematical relations to ergodicity and the Dry Ten Martini problem [27, 30, 29] present another direction for future work. It would for instance be worth exploring the case of several incommensurate frequencies, which are likely to result in enhanced stabilization mechanisms that in turn should lead to novel measurable signatures. These pursuits can also be followed in the context of the Dry Ten Martini problem without dualities or with hidden dualities [30, 29, 61], where localization phase transitions may be richer.
@@
In two-dimensional quasiperiodic systems, i.e. systems that are quasiperiodic along both dimensions, the presence of second Chern numbers and a richer set of topological invariants could lead to more examples of gap-independent topological protection. For example, in ultracold atomic systems, the scheme could be extended to the two-dimensional Aubry-André model or optical quasicrystals [49, 50, 51], where pumping could be realised along multiple dimensions. The recent advent of moiré van-der-Waals structures, in which pumping can be achieved upon applying periodic external gauge fields, provides an exciting solid-state platform in which these new mechanisms might both occur and help explain experimental observations.
@@
In summary, our results uncover a novel protection mechanism for Thouless pumping in quasiperiodic systems, highlighting an unexplored richness of topology in quasiperiodic systems and motivating a re-examination of topological protection in the absence of translation invariance. They furthermore enable novel experimental applications, notably in the preparation of quantum states with high Chern numbers.
@@ VI Acknowledgements
We cordially thank Joel E. Moore for insightful feedback and advice regarding the draft. We also thank Arjun Ashoka, Alexander Avdoshkin, Vir B. Bulchandani, Fabian Heidrich-Meisner, Matthew Henricks, Johannes Mitscherling, Suman Mondal, Lee Reeve, Ruben Verresen, Qijun Wu, and Maciej Zworski for many helpful discussions. R.-J.S. acknowledges funding from a New Investigator Award, EPSRC grant EP/W00187X/1, a EPSRC ERC underwrite grant EP/X025829/1, and a Royal Society exchange grant IES/R1/221060 as well as Trinity College, Cambridge. Work by D.S.B. was performed with support through a MURI project supported by the Air Force Office of Scientific Research (AFSOR) under grant number FA9550-22-1-027. U.S. and E.G. acknowledges support by the European Commission ERC Starting Grant QUASICRYSTAL, the EPSRC Grant (No. EP/R044627/1), and EPSRC Programme Grant DesOEQ (No. EP/P009565/1) and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via Research Unit FOR 2414 under project number 277974659. E.G. acknowledges support from the Cambridge Trust.
@@ VII Authors contributions
Appendix A Sign of gap labels from perturbation theory
@@
We showed in Section I how resonant pairs of lattice sites coupled by the effective $n$-th order tunneling element $J_{eff}^{(n)}$ give rise to pairs of energy gaps in the spectrum of the AAH chain at energies ±$\pm 2V\cos{(n\pi\beta)}$ ). According to the gap labeling theorem, each of these gaps is characterised by a Chern number, whose sign dictates the direction of the current arising under adiabatic pumping when all states below that gap are occupied.
@@
In the case $[n\beta]$ even, let us for instance fill up all states up to the energy $-2V\cos{(n\pi\beta)}$ ), see Fig. 8, which for the case discussed in Fig. 2 ($n=1$, $\beta>0.5$) is positive. For states below this gap, the pumping induces Landau-Zener hoppings in the negative direction, where each Landau-Zener transition transfers the states from the right-hand resonant neighbor to the left-hand one, see Fig. 8a. In this case we can ignore all other gaps, as they do not influence the overall current since their adjacent bands are either both filled or both empty. The pumping thus results in a quantized current in the negative direction and this gap is therefore associated with a negative Chern number. The same reasoning can be applied to the state filled up to $2V\cos{(n\pi\beta)}$ ), leading to a positive gap label.
@@
In the case $[n\beta]$ odd, the resonant neighbors are flipped, see Fig. 8b, and when filling all the states up to $-2V\cos{(n\pi\beta)}$ ), the adiabatic Landau-Zener transition transfers the states from the left- to the right-hand neighbor. In this case, the quantized current flows in the positive direction and is therefore associated to a positive Chern number.
@@
Therefore, for the sign of the gap labels to match the direction of the quantized current, one has to treat the even and odd cases of $[n\beta]$ separately, as done in Eq. 9.
@@
Figure 8: Dependence of the signs of the gap labels on the parity of $[n\beta]$: In the circle forming configuration space, the Landau-Zener transition for states below the gap always goes from left to right. However, depending on $[n\beta]$, the position of left- and right-hand neighbours on the circle can switch. Figures a and a show two different gaps at similar energies to illustrate the two cases.(a) When $[n\beta]$ is even, the associated Landau-Zener transition involves transport from the right-hand resonant neighbor to the left-hand one. The resulting quantized current flows in the negative direction and is therefore associated with a negative Chern number. (b) When $[n\beta]$ is odd, the situation is reversed, and the quantized current flows in the positive direction such that the associated Chern number is positive. Note that in contrast to Fig. 2, we are here considering states below the gap.
@@
Appendix B Estimation of critical disorder strengths for first-order $C=+1$ bands